OpenStudy (anonymous):

Evaluate. 8 * 3 + 6 ---------- -2 [6 / 2 - (9 - 1)]

OpenStudy (ajspeller):

Use the order of operations. Parentheses Exponents Multiplication - Division from Left to Right Addition - Subtraction from Left to Right

OpenStudy (anonymous):

8*3 = 24 +6 = 30 and 9 - 1 = 8 - 2 = 6 and 6/2 = 3*-2 = -6 / 30 I'm I right?

OpenStudy (daniellelovee):

so your final answer is -6/30??

OpenStudy (anonymous):

-5

OpenStudy (daniellelovee):

I got 3

OpenStudy (ajspeller):